OpenStudy (anonymous):

Assume that y varies inversely with x. If y = -21 when x = 4, find x when y = 7.

OpenStudy (zzr0ck3r):

now find x by letting y = 7

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

wait i thought y=-84

OpenStudy (anonymous):

@zzr0ck3r i dont get it

OpenStudy (zzr0ck3r):

ok we have the equation y=k/x and they give us a point (4,-21) and another rpoint (?,7) we need to figure out ? so first we don't even think about the x we need to figure out, because before we can do that we need to figure out what k is so we use the first point to figure out what k is y = k/x becomes -21 = k/4 so we solve for k k=-84 now we know our equation is like this y = -84/x (we now know what k is from our before work) no we go back to that other point (?,7) y = -84/x becomes 7 = -84/x solve for x to get the answer

OpenStudy (anonymous):

-1.3 ???

OpenStudy (anonymous):

wait no its -12

OpenStudy (zzr0ck3r):

7 = -84/x x = -84/7 x = -12

OpenStudy (anonymous):

YAY I GOT IT RIGHT thank youuuu

OpenStudy (anonymous):

i have one more! Assume that y varies inversely as x. If y = -8 when x = -9, find y when x = 4.

OpenStudy (zzr0ck3r):

do the exact same thing

OpenStudy (zzr0ck3r):

y=k/x use the first point (-9,-8) to solve for k then you x=4 and solve for y